# Phase dynamics and their role determining energy flux in hydrodynamic shell models

### v1

@@ I.1 Motivation
Why the instability assumption produces correct predictions is still an open question, particularly when considering the fact that the dynamics of a single triad represent a 2D three-component system which has more conserved quantities and different dynamics than that of 3D turbulence [51, 52]. Furthermore, none of the approaches described above predict how the fraction of energy going to large scales depends on the geophysically-relevant control parameter in anisotropic flows, let alone capture the apparent phase transition in the cascade which is observed [53, 54, 55, 56, 4, 31]. Our goal in this study is to take the next step towards an understanding of the energy cascade in turbulence (and its models) by introducing a new approach based on the dynamics of the complex phases of the Fourier amplitudes in spectral space. As a first step, we will uncover an approximate statistical description of these phases in shell models of hydrodynamic turbulence, which we will use to make new predictions about the cascade of conserved quantities.
@@ I.2 Background
In what follows, we make a crucial simplifying assumption on the dynamics of the triad phase for hydrodynamic shell models, resulting in an analytical approximation for the statistics of the triad phases. Given the new statistical description, we write, for the first time, an explicit expression for the average phase alignment and flux of conserved quantities for shell models as a function of the exponent of the kinetic energy spectrum as well as the second conserved quantity. These results are confirmed by numerical simulations of so-called phase-only shell models, where we fix the kinetic energy spectrum and evolve only the complex phases. Implications on the evolution of non-phase-only (i.e. ‘full’) shell models, past observations of shell models, as well as on an extension to the full Navier-Stokes equations are discussed in the conclusions.
@@ II.1 Hydrodynamic shell models
where $k_{n}=k_{0}\lambda^{n}$ is the wavenumber, $u_{n}(t)=u(k_{n},t)$ is the complex velocity amplitude for each mode (with $u^{*}_{n}$ its complex conjugate), $f_{n}$ is a forcing, $D_{n}$ is a dissipative mechanism, and $[a,b,c]$ are coefficients which determine the conserved quantities of the system. The index $n$ goes from 00 to $N-1$, where $N$ is the total number of shells in the system. The boundary conditions are $u_{-2}=u_{-1}=u_{N}=u_{N+1}=0$. The forcing and dissipative mechanism will remain unspecified for now.
Following standard practice, we choose $a=1$. Energy conservation requires that $a+b+c=0$, which we use to define $c=-a-b$. Our system then conserves the total energy,
@@
This leaves only one free parameter, $b$, which determines the second conserved quantity, which we will call $H$. For $-1<b<0$, $H$ is sign-indefinite, making it analogous to 3D turbulent flows, and is of the form,
@@
On the other hand, for $-2<b<-1$ the second conserved quantity is positive definite, making it analogous to 2D turbulent flows and is of the form,
@@
and not on the individual phases (with the exception of the forcing term). In fact, one can prove that the individual phases $\phi_{n}$ are uniformly distributed [76], yet an out-of-equilibrium solution requires that the triad phases not be uniformly distributed, making it the dynamical variable of interest. We therefore perform one last change of variables, making the primary variables $(\rho_{n},\theta_{n})$.
@@ II.2 Phase-only model
+(±$+\left(\pm\lambda^{2+\alpha-\gamma}-\lambda^{2+\alpha}-\lambda^{2+2\alpha}\right)\cos\left(\theta_{n+1}\right)$
@@
+(±$+\left(\pm\lambda^{2+\alpha-\gamma}-\lambda^{2+\alpha}-\lambda^{2+2\alpha}\right)\cos\left(\theta_{n+1}\right)$
@@
As is evident by Eq. (16), by fixing the shell magnitudes we explicitly ignore the contributions from their evolution and fluctuations on the dynamics of the triad phase. We also implicitly assume that correlations between fluctuating components of $\rho_{m}$ and $\theta_{n}$ are negligible for any pair of shells $m$ and $n$. Despite these strong assumptions, we found that the comparison of the statistics of $\theta_{n}$ between full 3D-like shell models and their phase-only counterpart (with the same spectral exponent) agree quite well, at least when considering the steady state probability distribution and $n$-independence. The two models will likely not agree when considering higher-order statistical measures, such as intermittency, but that is beyond the scope of this paper. As for 2D-like shell models, the power-law inertial range assumption and $n$-independence seem to break down in the full shell model. The reason for this is uncovered in our work on triad phase dynamics in the next section, and its implications will be discussed in the conclusion section.
@@ III Simplified triad phase dynamics: noisy phase oscillator
Our key simplifying assumption for this work is the following: we will treat the sum of all neighboring triads in Eq. (16) as a single noise variable $\xi_{n}$, in other words $\sum_{m\neq n}K_{mn}\cos(\theta_{m})\mapsto\xi_{n}$. In doing so, we implicitly neglect any statistical dependence between pairs of triad phases $\theta_{n},\theta_{m}$ for all $n\neq m$, and invoke the central limit theorem to ensure that $\xi_{n}$ is a Gaussian variable. Namely, our key assumption neglects the phenomenon of triad-phase synchronization, whereby triad phases corresponding to different scales are correlated in time so as to enhance the flux. Such a phenomenon is observed in the 1D Burgers equation [59] and in shell models [74, 75], and is directly related to extreme events and intermittency. Statistical independence of the triad phases was first proposed as a simplifying assumption in a cascade model of turbulence, with a focus on intermittency [76]. However, shell model simulations show that this is not entirely true. Indeed, this cannot hold if intermittency is to be present in the system, since the only constant flux solution for independent triad phases, as mentioned in the previous section, is the self-similar Kolmogorov solution [78]. Although our key assumption may not precisely hold, we show here that it does indeed provide a first-order description of the steady state statistics of $\theta_{n}$ and its implications on the flux, which are our main goals for this work. In summary, our results should be valid in a regime where extreme events and intermittency do not contribute significantly to the statistics of the triad phases.
@@
Equation (18) is that of a noisy phase oscillator, also known as the Adler equation (with an intrinsic frequency equal to zero). The deterministic dynamics of the triad phase are thus solely determined by the so-called ‘self-interaction term’, $\mathcal{K}\cos(\theta_{n})$. The simplicity of the deterministic Adler equation gives us direct access to the fixed points and their stability. The two fixed points of Eq. (18) are ±$\pm\pi/2$, with $\pi/2$ being stable for $\mathcal{K}>0$ whereas $-\pi/2$ is stable for $\mathcal{K}<0$. This already hints at implications of the sign of $\mathcal{K}$ on the sign of the alignment strength, which is $\sin(\theta)=\pm 1$ for the deterministic fixed point values of $\theta=\pm\pi/2$, as well as the sign of the flux.
Before moving on to the full implications of the nondeterministic dynamics of our model for the triad phase evolution, we investigate the validity of our assumptions using a pair of phase-only simulations in the 3D- and 2D-like regimes. Both runs have $N=28$ modes, $\lambda=1.5$, $k_{0}=\lambda^{-14}$, $\rho_{0}=10^{-0.5}k_{0}^{\alpha/2}$, $a=1$, and were initialized with random phases. The 3D-like run has a value of $\alpha=-2/3$ and $\gamma=1$ (so that $b=-1/3$), whereas the 2D-like run has a value of $\alpha=-2$ and $\gamma=0.855$ (so that $b=-1.707$). Equation (11) was integrated in time using a fourth order Runge-Kutta method with time step $\Delta t=10^{-2}$. We test our assumptions by outputting the supposed ‘noise’ variable $\widetilde{\xi}_{n}=\sum_{m\neq n}K_{mn}\cos(\theta_{m})$ for a few modes $n=\{N/4,N/3,N/2\}$. Note that for our analytical considerations $\xi_{n}$ is a true Gaussian random variable, whereas $\widetilde{\xi}_{n}$ is deterministic, albeit chaotic. Our goal is to see how close $\widetilde{\xi}_{n}$ comes to being a true Gaussian random variable like $\xi_{n}$, and to investigate its properties for modeling purposes.
@@
Figure 1: Statistics of $\widetilde{\xi}_{n}$, the ‘noise’ variable measured from phase-only shell model runs of a 2D-like case (top row, (a) and (b)) and a 3D-like case (bottom row, (c) and (d)). The PDFs both show an approximately Gaussian shape, with a maximum at zero and a fall-off towards zero. Both autocorrelation functions have time-scales of order one. That of the 2D-like case, (b), shows a decaying and oscillating autocorrelation function, hinting at colored harmonic noise. Each plot shows statistics for $n=N/4$ (blue), $n=N/3$ (orange), and $n=N/2$ (green). The overlap of these curves demonstrates the $n$-independent statistics of the re-scaled variables.
@@
Under a few reasonable assumptions, it is possible to compute the steady-state statistics of Eq. (18) taking $\xi_{n}$ to be colored harmonic noise (see the Appendix). Indeed, this gives the best fits to the statistics of $\theta_{n}$. Although more accurate, we found that the complexity of the resulting expressions for the statistics limited our ability to interpret and understand the results, and made the predictions of quantities of interest more difficult to compute. Considering that our goal is to give a first-order description, we choose instead to model our noise using the simplest possible choice: white noise, such that $\overline{\xi_{n}(t)\xi_{n}(t+s)}=D_{\mathrm{eff}}\delta(s),$ where $\overline{\cdot}$ denotes an ensemble average and $D_{\mathrm{eff}}$ is the resulting ‘effective’ noise amplitude which best matches the data. Modeling the noise as white, in turn, reduces the undetermined free parameters in the model from three (for the colored harmonic noise) to one (for the white noise). This reduces the sensitivity of the fits to be performed in subsequent sections: whereas the colored harmonic noise fits are non-unique and depend strongly on initial guesses, the white noise fits for $D_{\mathrm{eff}}$ are robust. Even though $\widetilde{\xi}_{n}$ clearly has a similar time-scale as $\theta_{n}$, modeling $\xi_{n}$ as white noise will turn out to be sufficient to describe the steady-state statistics of $\theta_{n}$ to first order. In what follows we drop subscripts $n$, since the steady-state statistics do not depend on the mode $n$.
@@
Figure 2: A plot of the predicted PDF of $\theta$, Eq. (20), based on our model for the triad phase dynamics, Eq. (18), assuming white noise. Multiple values of $\mathcal{K}/D_{\mathrm{eff}}$ are shown, with red color corresponding to positive $\mathcal{K}/D_{\mathrm{eff}}$, blue to negative $\mathcal{K}/D_{\mathrm{eff}}$, and black to $\mathcal{K}/D_{\mathrm{eff}}=0$. Notice how the position of the maximum value of the PDF, ±$\pm\pi/2$, changes sign depending on the sign of $\mathcal{K}/D_{\mathrm{eff}}$.
@@ IV Comparison with shell model simulations
We performed a total of 600 phase-only shell model runs (312 for the 2D-like runs and 288 for the 3D-like runs), varying the spectral slope $\alpha$ and the exponent $\gamma$ of $k_{n}$ in the second conserved quantity $H$. For both sets of runs, we varied the spectral slope between $\alpha=-2$ and $\alpha=0$, with increments $\Delta\alpha=0.087$ resulting in 24 different values of $\alpha$. For the 2D-like runs, the exponent of $k_{n}$ in $H$ was varied between $\gamma=0.13$ (corresponding to $b=-1.95$) and $\gamma=3.42$ (corresponding to $b=-1.25$), with a constant increment in $b$, $\Delta b=0.05$, resulting in 13 different values of $\gamma$. For the 3D-like runs, $\gamma$ was varied between $\gamma=0.26$ (corresponding to $b=-0.1$) and $\gamma=2.97$ (corresponding to $b=-0.7$), with an approximately constant increment in $\gamma$, $\Delta\gamma\approx 0.25$, resulting in 12 values of $\gamma$.
@@ IV.1 Positive-definite $H$
We first focus on the 2D-like runs, whose second conserved quantity $H$ is positive-definite. Figure 3(a) shows PDFs of the triad phase for five different runs, chosen to demonstrate the variety of behavior found, including different alignment strengths, PDF maxima at ±$\pm\pi/2$, and an unsynchronized uniform distribution.
@@
We remind the reader that the measurement of what we call the energy flux, Eq. (17), is really a measure of the energy flux that would occur in the full shell model, given a power law inertial range with spectral exponent $\alpha$ and $n$-independent triad phase statistics. Therefore, our results do not suggest that full models with positive-definite $H$ will have a sustained forward energy cascade. Instead, they reveal which inertial range power law exponents support an inverse energy cascade. Take, for example, the Kolmogorov self-similar solution of $\alpha=-2/3$ at $\gamma=0.85$. For an inverse cascading system, we must have negative alignment, meaning $\theta_{n}=-\pi/2$. With these values of $\alpha$ and $\gamma$, this solution sits above the $\mathcal{K}=0$ line of $\alpha_{c,2}$ (Fig. 4(a)), and our model for the phase dynamics suggests that this fixed point is unstable: instead, it is the $\theta_{n}=\pi/2$ fixed point that is stable. Therefore, the Kolmogorov self-similar solution for an inverse cascade range is not stable and the system must break either the assumption of $n$-independent triad phase statistics or of having $\alpha=-2/3$. Indeed, a full model simulation of the system at $\gamma=0.85$ (not shown) reveals that both are true: the system forms an inverse cascade range where $\langle\sin(\theta_{n})\rangle$ varies with $n$ and which has an exponent $\alpha\approx-1.14$. This value of $\alpha$ now lies between $\alpha_{c,2}$ and $\alpha_{c,3}$ in the range where $\theta_{n}=-\pi/2$ is a stable fixed point of the model dynamics. The breakdown of the $n$-independence of triad phase statistics is necessary for the presence of an energy flux that is constant in $n$ (Section II.2).

### v2

@@ I.1 Motivation
Why the instability assumption produces correct predictions is still an open question, particularly when considering the fact that the dynamics of a single triad represent a 2D three-component system which has more conserved quantities and different dynamics than that of 3D turbulence [61, 13]. Furthermore, none of the approaches described above predict how the fraction of energy going to large scales depends on the geophysically-relevant control parameter in anisotropic flows, let alone capture the apparent phase transition in the cascade which is observed [72, 71, 11, 65, 1, 78, 79]. Our goal in this study is to take the next step towards an understanding of the energy cascade in turbulence (and its models) by introducing a new approach based on the dynamics of the complex phases of the Fourier amplitudes in spectral space. As a first step, we will uncover an approximate statistical description of these phases in shell models of hydrodynamic turbulence, which we will use to make new predictions about the cascade of conserved quantities.
@@ I.2 Background
In what follows, we make a crucial simplifying assumption on the dynamics of the triad phase for hydrodynamic shell models, resulting in an analytical approximation for the statistics of the triad phases. Given the new statistical description, we write, for the first time, an explicit expression for the average phase alignment and flux of conserved quantities for shell models as a function of the exponent of the kinetic energy spectrum as well as the second conserved quantity. These results are confirmed by numerical simulations of so-called phase-only shell models, where we fix the kinetic energy spectrum and evolve only the complex phases. Implications on the evolution of non-phase-only (i.e. ‘full’) shell models, past observations of shell models, as well as on an extension to the full Navier-Stokes equations are discussed in the conclusions and in Appendix B.
@@ II.1 Hydrodynamic shell models
where $k_{n}=k_{0}\lambda^{n}$ is the wavenumber, $u_{n}(t)=u(k_{n},t)$ is the complex velocity amplitude for each mode (with $u^{*}_{n}$ its complex conjugate), $f_{n}$ is a forcing, $D_{n}$ is a dissipative mechanism, and $[a,b,c]$ are coefficients which determine the conserved quantities of the system. The index $n$ goes from 0 to N1N-1, where $N$ is the total number of shells in the system. The boundary conditions are $u_{-2}=u_{-1}=u_{N}=u_{N+1}=0$. The forcing and dissipative mechanism will remain unspecified for now.
Following standard practice, we choose $a=1$. Energy conservation requires that $a+b+c=0$, which we use to define c=abc=-a-b. Our system then conserves the total energy,
@@
This leaves only one free parameter, $b$, which determines the second conserved quantity, which we will call $H$. For 1<b<0-1<b<0, $H$ is sign-indefinite, making it analogous to 3D turbulent flows, and is of the form,
@@
On the other hand, for 2<b<1-2<b<-1 the second conserved quantity is positive definite, making it analogous to 2D turbulent flows and is of the form,
@@
and not on the individual phases (with the exception of the forcing term). In fact, one can prove that the individual phases $\phi_{n}$ are uniformly distributed [12], yet an out-of-equilibrium solution requires that the triad phases not be uniformly distributed, making it the dynamical variable of interest. We therefore perform one last change of variables, making the primary variables $(\rho_{n},\theta_{n})$. Note that, although the triad phase $\theta_{n}$ has a single index $n$, it comprises phases from three different modes: $n,n+1,$ and $n+2$.The single index is due to the fact that in shell models, by construction, any triad consists of three “contiguous” neighbors, and so one can choose for example the lowest shell index in the triad to “represent” the triad.
@@ II.2 Phase-only model
$+\left(\pm\lambda^{2+\alpha-\gamma}-\lambda^{2+\alpha}-\lambda^{2+2\alpha}\right)\cos\left(\theta_{n+1}\right)$
@@
$+\left(\pm\lambda^{2+\alpha-\gamma}-\lambda^{2+\alpha}-\lambda^{2+2\alpha}\right)\cos\left(\theta_{n+1}\right)$
@@
As is evident by Eq. (16), by fixing the shell magnitudes we explicitly ignore the contributions from their evolution and fluctuations on the dynamics of the triad phase. We also implicitly assume that correlations between fluctuating components of $\rho_{m}$ and $\theta_{n}$ are negligible for any pair of shells $m$ and $n$. Despite these strong assumptions, we found that the comparison of the statistics of $\theta_{n}$ between full 3D-like shell models and their phase-only counterpart (with the same spectral exponent) agree quite well, at least when considering the steady state probability distribution and $n$-independence. The two models will likely not agree when considering higher-order statistical measures, such as intermittency, but that is beyond the scope of this paper. As for 2D-like shell models, the power-law inertial range assumption and $n$-independence seem to break down in the full shell model (Appendix B). The reason for this is uncovered in our work on triad phase dynamics in the next section, and its implications will be discussed in the conclusion section.
@@ III Simplified triad phase dynamics: noisy phase oscillator
Our key simplifying assumption for this work is the following: we will treat the sum of all neighboring triads in Eq. (16) as a single noise variable $\xi_{n}$, in other words $\sum_{m\neq n}K_{mn}\cos(\theta_{m})\mapsto\xi_{n}$. In doing so, we implicitly neglect any statistical dependence between pairs of triad phases $\theta_{n},\theta_{m}$ for all $n\neq m$, and invoke the central limit theorem to ensure that $\xi_{n}$ is a Gaussian variable. Namely, our key assumption neglects the phenomenon of triad-phase synchronization, whereby triad phases corresponding to different scales are correlated in time so as to enhance the flux. Such a phenomenon is observed in the 1D Burgers equation [63] and in shell models [22, 56], and is directly related to extreme events and intermittency. Statistical independence of the triad phases was first proposed as a simplifying assumption in a cascade model of turbulence, with a focus on intermittency [12]. However, shell model simulations show that this is not entirely true. Indeed, this cannot hold if intermittency is to be present in the system, since the only constant flux solution for independent triad phases, as mentioned in the previous section, is the self-similar Kolmogorov solution [31]. Although our key assumption may not precisely hold, we show here that correlations between neighboring triad phases are weak enough that our model does indeed provide a sufficiently accurate description of the steady state statistics of $\theta_{n}$ and its implications on the flux, which are our main goals for this work. In summary, our results should be valid in a regime where extreme events and intermittency do not contribute significantly to the statistics of the triad phases.
@@
Equation (18) is that of a noisy phase oscillator, also known as the Adler equation (with an intrinsic frequency equal to zero). The deterministic dynamics of the triad phase are thus solely determined by the so-called ‘self-interaction term’, $\mathcal{K}\cos(\theta_{n})$. The simplicity of the deterministic Adler equation gives us direct access to the fixed points and their stability. The two fixed points of Eq. (18) are $\pm\pi/2$, with $\pi/2$ being stable for $\mathcal{K}>0$ whereas $\pi/2$ is stable for $\mathcal{K}<0$. This already hints at implications of the sign of $\mathcal{K}$ on the sign of the alignment strength, which is $\sin(\theta)=\pm 1$ for the deterministic fixed point values of $\theta=\pm\pi/2$, as well as the sign of the flux.
Before moving on to the full implications of the nondeterministic dynamics of our model for the triad phase evolution, we investigate the validity of our assumptions using a pair of phase-only simulations in the 3D- and 2D-like regimes. Both runs have $N=28$ modes, $\lambda=1.5$, $k_{0}=\lambda^{-14}$, $\rho_{0}=10^{-0.5}k_{0}^{\alpha/2}$, $a=1$, and were initialized with random phases. The 3D-like run has a value of $\alpha=-2/3$ and $\gamma=1$ (so that b=1/3b=-1/3), whereas the 2D-like run has a value of $\alpha=-2$ and $\gamma=0.855$ (so that b=1.707b=-1.707). Equation (11) was integrated in time using a fourth order Runge-Kutta method with time step $\Delta t=10^{-2}$. We test our assumptions by outputting the supposed ‘noise’ variable $\widetilde{\xi}_{n}=\sum_{m\neq n}K_{mn}\cos(\theta_{m})$ for a few modes $n=\{N/4,N/3,N/2\}$. Note that for our analytical considerations $\xi_{n}$ is a true Gaussian random variable, whereas $\widetilde{\xi}_{n}$ is deterministic, albeit chaotic. Our goal is to see how close $\widetilde{\xi}_{n}$ comes to being a true Gaussian random variable like $\xi_{n}$, and to investigate its properties for modeling purposes.
@@
Figure 1: Statistics of $\widetilde{\xi}_{n}$, the ‘noise’ variable measured from phase-only shell model runs of a 2D-like case (top row, (a) and (b)) and a 3D-like case (bottom row, (c) and (d)). The PDFs both show an approximately Gaussian shape(dashed red line), with a maximum at zero and a fall-off towards zero. Both autocorrelation functions have time-scales of order one. That of the 2D-like case, (b), shows a decaying and oscillating autocorrelation function, hinting at colored harmonic noise. Each plot shows statistics for $n=N/4$ (blue), $n=N/3$ (orange), and $n=N/2$ (green). The overlap of these curves demonstrates the $n$-independent statistics of the re-scaled variables.
@@
Under a few reasonable assumptions, it is possible to compute the steady-state statistics of Eq. (18) taking $\xi_{n}$ to be colored harmonic noise (see Appendix A). Indeed, this gives the best fits to the statistics of $\theta_{n}$. Although more accurate, we found that the complexity of the resulting expressions for the statistics limited our ability to interpret and understand the results, and made the predictions of quantities of interest more difficult to compute. Considering that our goal is to give a first-order description, we choose instead to model our noise using the simplest possible choice: white noise, such that $\overline{\xi_{n}(t)\xi_{n}(t+s)}=D_{\mathrm{eff}}\delta(s),$ where $\overline{\cdot}$ denotes an ensemble average and $D_{\mathrm{eff}}$ is the resulting ‘effective’ noise amplitude which best matches the data. Modeling the noise as white, in turn, reduces the undetermined free parameters in the model from three (for the colored harmonic noise) to one (for the white noise). This reduces the sensitivity of the fits to be performed in subsequent sections: whereas the colored harmonic noise fits are non-unique and depend strongly on initial guesses, the white noise fits for $D_{\mathrm{eff}}$ are robust. Even though $\widetilde{\xi}_{n}$ clearly has a similar time-scale as $\theta_{n}$, modeling $\xi_{n}$ as white noise will turn out to be sufficient to describe the steady-state statistics of $\theta_{n}$ to first order. In what follows we drop subscripts $n$, since the steady-state statistics do not depend on the mode $n$.
@@
Figure 2: A plot of the predicted PDF of $\theta$, Eq. (20), based on our model for the triad phase dynamics, Eq. (18), assuming white noise. Multiple values of $\mathcal{K}/D_{\mathrm{eff}}$ are shown, with red color corresponding to positive $\mathcal{K}/D_{\mathrm{eff}}$, blue to negative $\mathcal{K}/D_{\mathrm{eff}}$, and black to $\mathcal{K}/D_{\mathrm{eff}}=0$. Notice how the position of the maximum value of the PDF, $\pm\pi/2$, changes sign depending on the sign of $\mathcal{K}/D_{\mathrm{eff}}$.
@@ IV Comparison with shell model simulations
We performed a total of 600 phase-only shell model runs (312 for the 2D-like runs and 288 for the 3D-like runs), varying the spectral slope $\alpha$ and the exponent $\gamma$ of $k_{n}$ in the second conserved quantity $H$. For both sets of runs, we varied the spectral slope between $\alpha=-2$ and $\alpha=0$, with increments $\Delta\alpha=0.087$ resulting in 24 different values of $\alpha$. For the 2D-like runs, the exponent of $k_{n}$ in $H$ was varied between $\gamma=0.13$ (corresponding to b=1.95b=-1.95) and $\gamma=3.42$ (corresponding to b=1.25b=-1.25), with a constant increment in $b$, $\Delta b=0.05$, resulting in 13 different values of $\gamma$. For the 3D-like runs, $\gamma$ was varied between $\gamma=0.26$ (corresponding to b=0.1b=-0.1) and $\gamma=2.97$ (corresponding to b=0.7b=-0.7), with an approximately constant increment in $\gamma$, $\Delta\gamma\approx 0.25$, resulting in 12 values of $\gamma$.
@@ IV.1 Positive-definite $H$
We first focus on the 2D-like runs, whose second conserved quantity $H$ is positive-definite. Figure 3(a) shows PDFs of the triad phase for five different runs, chosen to demonstrate the variety of behavior found, including different alignment strengths, PDF maxima at $\pm\pi/2$, and an unsynchronized uniform distribution.
@@
We remind the reader that the measurement of what we call the energy flux, Eq. (17), is really a measure of the energy flux that would occur in the full shell model, given a power law inertial range with spectral exponent $\alpha$ and $n$-independent triad phase statistics. Therefore, our results do not suggest that full models with positive-definite $H$ will have a sustained forward energy cascade. Instead, they reveal which inertial range power law exponents support an inverse energy cascade. Take, for example, the Kolmogorov self-similar solution of $\alpha=-2/3$ at $\gamma=1.26$. For an inverse cascading system, we must have negative alignment, meaning $\theta_{n}=-\pi/2$. With these values of $\alpha$ and $\gamma$, this solution sits above the $\mathcal{K}=0$ line of $\alpha_{c,2}$ (Fig. 4(a)), and our model for the phase dynamics suggests that this fixed point is unstable: instead, it is the $\theta_{n}=\pi/2$ fixed point that is stable. Therefore, the Kolmogorov self-similar solution for an inverse cascade range is not stable and the system must break either the assumption of $n$-independent triad phase statistics or the assumption of having $\alpha=-2/3$, or both. Indeed, a full model simulation of the system at $\gamma=1.26$ (Fig. 6, Appendix B) reveals that both assumptions are broken: the system forms an inverse cascade range where $\langle\sin(\theta_{n})\rangle$ varies with $n$ and which has an exponent $\alpha\approx-1.32$. This value of $\alpha$ now lies between $\alpha_{c,2}$ and $\alpha_{c,3}$ in the range where $\theta_{n}=-\pi/2$ is a stable fixed point of the model dynamics. In fact, Figure 7 in Appendix B shows that the same is true for all full runs: the final exponents of the inverse cascading inertial range all lie within the regions where negative flux is possible, based on our model’s prediction. This is rather surprising, given the fact that our zero-flux lines are based on the assumption that the statistics are $n$-independent, yet all of the full runs have clearly $n$-dependent phase alignment. The breakdown of the $n$-independence of triad phase statistics is necessary for the presence of an energy flux that is constant in $n$ (Section II.2, Fig. 6). Many possible energy spectra and triad phase distributions can be consistent with a particular constant flux solution. As we have seen above, considering the triad phase dynamics therefore acts to select the possible dynamical stable configuration that would be found in a full shell model simulation.